OpenStudy (anonymous):

How do you find the arc length? Explain the two methods learned in this section.

OpenStudy (anonymous):

forget the "in this section" part. i just need two ways

hero (hero):

Two methods of finding arc length: \[s = r\theta\] s = arc length r = radius \[\frac{x}{360} = \frac{y}{2\pi r}\] x = angle that spans arc length y = arc length

OpenStudy (anonymous):

you really ARE MY HERO!!!

OpenStudy (anonymous):

Um, i do have a small question

OpenStudy (anonymous):

ho do i type that? i cant copy and paste that

hero (hero):

How do you type it on another webpage...or how do you type it here?

hero (hero):

If you're trying to type it on another webpage...I can't help you.

OpenStudy (anonymous):

how would i type it into a text box... yah like, on another page

hero (hero):

You'll have to find a theta symbol online that you can copy.

hero (hero):

Hang on

hero (hero):

θ

hero (hero):

For fractions just use the slash symbol /

OpenStudy (anonymous):

thanks. and is that an 8 in frront of "=θr"

hero (hero):

an "es" S

hero (hero):

I'm pretty sure you've seen an "es" before. It is a letter of the alphabet.

hero (hero):

It comes after r and before t

hero (hero):

It's just a variable.

OpenStudy (anonymous):

s lol. ur the bomb man

hero (hero):

If you've never seen that formula before....

hero (hero):

maybe it's not what your teacher taught?

hero (hero):

By the way, when did school start?